a manufacture has 15 years, 8.5 mortgage for 100,000 birr on a building payment are made monthly.

a. find the month payment

b. how much interest will be paid


1
Expert's answer
2022-03-03T12:34:57-0500

Loan amount= 100,000 birr

Monthly rate = 8.5/12= 0.71%

Number of periods (n) = 15 years* 12 months = 180 months


a)

Monthly payment "=\frac {100,000} {{1-\frac {1}{1+ 0.0071}^{180}}}"


"=\frac {100,000} {{1-\frac {1}{1.0071}^{180}}}"

"=\frac {100,000}{{\frac{1-28} {0.0071}}}"

"=\frac {100,000}{101.43}"

"= 985.91 \space birr"

b)

"=\frac{985.91}{101.91}"


"= 9.72\%"